# AnomalyDINO: Boosting Patch-based Few-shot Anomaly Detection with DINOv2

### v1

@@ 3.1 Anomaly Detection with Nearest Patch Distances
Let us consider the image space $\mathcal{X}$. We assume to have access to a suitable feature extractor $f:\mathcal{X}\rightarrow\mathcal{F}^{n}$ that maps each image $\mathbf{x}$ to a tuple of patch-features $f(\mathbf{x})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$, where $\mathcal{F}$ denotes the feature space. Note that $n$ depends on the image resolution and the patch size. Given $k\geq 1$ nominal reference sample(s) $X_{\text{ref}}:=\bigl{\{}\mathbf{x}^{(i)}\mid\,i\in\{1,\dots,k\}\bigr{\}}$ ($k$-shot setting), we collect the nominal patch features and store them in a memory bank
@@
$\mathcal{M}:=\bigcup_{\mathbf{x}^{(i)}\in X_{\operatorname{ref}}}\bigl{\{}\mathbf{p}_{j}^{(i)}\;|\;f(\mathbf{x}^{(i)})=(\mathbf{p}_{1}^{(i)},\dots,\mathbf{p}_{n}^{(i)}),\,j\in\{1,\dots,n\}\bigr{\}}\enspace.$ (1)
@@
To score a test sample $\mathbf{x}_{\mathrm{test}}$, we collect the extracted patch representations $f(\mathbf{x}_{\mathrm{test}})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$ and then check how well they comply with our memory bank $\mathcal{M}$. To do so, we leverage a nearest-neighbor approach to find the distance to the closest reference patch for a given test patch $\mathbf{p}\in\mathcal{F}$
@@
Throughout this paper, we define $q$ as the average distance of the 1% most anomalous patches, i.e, $q(\mathcal{D}):=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ with $H_{0.01}(\mathcal{D})$ containing the 1% highest values in $\mathcal{D}$. The statistic $q$ can be understood as an empirical estimate of the tail value at risk for the 99% quantile (e.g., McNeil, 1999) and turns out suitable for a wide range of settings because it balances two desirable properties: First, we want $s(\mathbf{x}_{\mathrm{test}})$ to depend on the highest patch distances, as they may provide the strongest anomaly signal. Similarly, we want a certain degree of robustness against a singular high patch distance (in particular in few-shot scenarios where $\mathcal{M}$ is sparsely populated). However, one could also replace $q$ with a different statistic to cater to special cases. If anomalies are expected to cover larger parts of the image, for example, a high percentile of the patch distances might be a suitable choice. If, on the contrary, anomalies may occur only locally such that only very few patches/or a single patch might be affected, the score $q$ should be sensitive to the highest patch distances. Frequently, the maximum pixel-wise anomaly score is considered. Such an anomaly score on pixel-level is usually obtained by upsampling to full image resolution and applying some smoothing operation.
Following Roth et al. (2022), we utilize bilinear upsampling and Gaussian smoothing (with $\sigma=4.0$) to turn the patch distances into pixel-level anomaly scores for localization of potential defects. Examples of the resulting anomaly maps are visualized in Figure 1 and Appendix A (Figures 3 and 4).
@@ 3.2 Enriching the Memory Bank & Filtering Relevant Patches
Access to only a single/few normal sample(s) poses (at least) two challenges. First, all relevant information about the concept of normality must be extracted from the single/few available nominal sample(s). However, it is expected that the variability at test time is (considerably) higher than in $X_{\mathrm{ref}}$. Secondly, one must determine a suitable preprocessing for reference and test samples, solely based on $X_{\mathrm{ref}}$. To address low variations of nominal patch features, we apply simple augmentations in the form of rotations, following the insights of Xie et al. (2023). In addition, to ensure that $\mathcal{M}$ is of high quality, we consider masking the object of interest as background noise could poison $\mathcal{M}$. Note, that a suitable preprocessing, i.e., whether to apply masking or rotations, should depend on the object of interest (see, e.g., Figures 7 and 10, and Section C.1 for further discussion).
Figure 2: Masking test on MVTec-AD. The dashed red regions mark areas that belong to the object of interest but are falsely predicted to be background. See App. C.1 for the outcomes per category.
@@
To minimize the overhead of the proposed method, we utilize DINOv2 also for masking, i.e., delineating the primary object(s) in an image from its background. This is achieved by thresholding the first PCA component of the patch features (Oquab et al., 2024). We observe empirically that this procedure sometimes produces erroneous masks. Frequently, such failure cases occur for close-up shots, where the objects of interest account for $\gtrsim 50\%$ of patches, see Figure 2. To resolve that issue we test whether the PCA-based mask correctly captures the object in the first reference sample and apply masking accordingly. In general, we do not mask textures (e.g., wood or tile in MVTec-AD).
@@
Rotating the reference sample may improve the detection performance by enriching the memory bank, but can also be detrimental in cases where rotations of (parts of the) object of interest can be considered anomalies themselves. We consider two different settings.
In the ‘agnostic’ case, which we focus on in the main paper, we always augment the reference sample with rotations. We also consider the case where we know about the samples’ potential rotations (‘informed’). The ‘informed’ case is sensible as the data collection process can usually be controlled in an industrial/medical setting (or test images can be aligned), and may lead to significantly lower inference time. See Section C.1 for the effect of masking and rotation, and a comparison of ‘informed’ vs. ‘agnostic’.
@@ 4 Experiments
Using DINOv2111Code and model weights of DINOv2 (Oquab et al., 2024) are available under Apache 2.0 license. as backbone allows us to choose from the distillation sizes, which range from small (ViT-S, $21\times 10^{6}$ parameters) to giant (ViT-G, $1.1\times 10^{9}$ parameters). To prioritize low latency, we take the smallest model (S) as our default. As discussed above, the ‘agnostic’ preprocessing is our default.
@@
For the experiments, we consider MVTec-AD [^2] The MVTec-AD dataset (Bergmann et al., 2021) is available under the CC BY-NC-SA 4.0 license. and VisA, [^3] The VisA dataset (Zou et al., 2022) is released under the CC BY 4.0 license. two datasets with high-resolution images for industrial anomaly detection. MVTec-AD consists of fifteen categories, both objects and textures, with a small amount of clean training samples (between 60 and 391), and up to eight anomaly types per category. VisA has twelve categories with up to 1006 normal samples per category and diverse types of anomalies condensed into one anomaly class ‘bad’.
@@
We assess the image-level detection and the pixel-level segmentation performance with three metrics each (similar to, e.g., Jeong et al., 2023; Chen et al., 2023; Li et al., 2024b). For evaluating the detection performance, we measure the area under the receiver-operator curve (AUROC), the F1-score at the optimal threshold (F1-max), and the average precision (AP) using the respective image-level anomaly scores. We quantify the anomaly segmentation performance using the AUROC, F1-max, and the per-region overlap (PRO, Bergmann et al., 2021) of the segmentation using the pixel-wise anomaly scores. Note in this context that due to a high imbalance of nominal and anomalous pixels—for MVTec-AD we have 97.26% nominal pixel, for VisA even 99.45% (Li et al., 2024b)—it is not recommended to assess performance solely on segmentation AUROC (Rafiei et al., 2023). We repeat each experiment three times and report mean and standard deviation. [^4] The randomness stems from the choice of reference samples $X_{\mathrm{ref}}$. For straightforward reproducibility, we simply set $X_{\mathrm{ref}}$ to the first $k$, second $k$ and third $k$ train samples for the three different runs, respectively.
@@
Setting Method Detection Segmentation
@@
AnomalyDINO 96.6$\pm$0.5 96.0$\pm$0.2 98.2$\pm$0.3 96.3$\pm$0.1 58.2$\pm$0.8 91.9$\pm$0.1
2-shot SPADE 81.0$\pm$2.0 90.3$\pm$0.8 90.6$\pm$0.8 91.2$\pm$0.4 42.4$\pm$1.0 83.9$\pm$0.7
@@
4-shot SPADE 84.8$\pm$2.5 91.5$\pm$0.9 92.5$\pm$1.2 92.7$\pm$0.3 46.2$\pm$1.3 87.0$\pm$0.5
@@
AnomalyDINO 98.4$\pm$0.1 97.8$\pm$0.2 99.3$\pm$0.0 97.1$\pm$0.1 60.2$\pm$0.1 93.0$\pm$0.1
Baselines
We compare AnomalyDINO with a range of modern zero- and few-shot AD models, e.g., SPADE (Cohen and Hoshen, 2020), PaDiM (Defard et al., 2020), PatchCode (Roth et al., 2022), WinCLIP+ (Jeong et al., 2023), and APRIL-GAN (Chen et al., 2023). It is important to note that ACR (Li et al., 2024a) and MuSc (Li et al., 2024b) consider the batched zero-shot setting (see Table 1), thus, covering a different setting than AnomalyDINO. We adapt AnomalyDINO to this setting in Appendix D. Moreover, APRIL-GAN and AnomalyCLIP, require training on a related dataset, which is in contrast to our proposed training-free method. In Tables 2 and 3, results reported by Jeong et al. (2023) are indicated by , the result of the WinCLIP re-implementation by Kwak et al. (2024) (where J. Jeong is also coauthor) by . All other results are taken from the original publication. For ADP (Kwak et al., 2024) we report the (usually slightly better) variant ADP (with access to the class label). For GraphCore no standard deviations are reported (Xie et al., 2023).
4.1 Few-Shot Anomaly Detection and Anomaly Segmentation
Results
We summarize the results for few-shot anomaly detection on MVTec-AD and VisA in Table 2 and Table 3, respectively. Regarding MVTec-AD, our method achieves state-of-the-art $k$-shot detection performance across all $k\in\{1,2,4,8,16\}$ in every reported metric, outperforming even approaches that require training on an additional data set (e.g., ADP and APRIL-GAN).
@@ Results
Setting Method Detection Segmentation
@@
[…]
AnomalyDINO reduces the gap between few-shot and full-shot methods in anomaly detection significantly, and it demonstrates the strongest performance in anomaly localization. Furthermore, we see a clear improvement in the method’s results as the quantity of samples increases.
In terms of anomaly detection in the VisA benchmark (Table 3), APRIL-GAN demonstrates the strongest performance for $k\in\{1,2,4\}$. Nonetheless, AnomalyDINO consistently achieves second-best results, comparable results with APRIL-GAN in the 8-shot setting, and sets new state-of-the-art for $k=16$. This can be attributed to AnomalyDINO’s ability to benefit more from a richer memory bank $\mathcal{M}$ than APRIL-GAN. We hypothesize that meta-learning exerts a greater influence on APRIL-GAN (i.e., training on MVTec-AD, when testing on VisA, and vice-versa) compared to learning from the nominal features of the given reference samples.Note also that, AnomalyDINO outperforms all other training-free approaches.
Table 4: Detection results for other settings. All results are AUROC values in (%).
@@
AnomalyDINO 93.3 89.6
Comparing the segmentation performance on VisA, Table 3 reveals a mixed picture: WinCLIP+ shows strong localization performance in terms of pixel-level AUROC, and APRIL-GAN achieves the highest scores in PRO. AnomalyDINO consistently achieves results in per-region-overlap that are second-best compared to all other, and the best among the training-free methods. In addition, AnomalyDINO shows the best segmentation performance when measured in F1-max, a metric that accounts for the strong imbalance of nominal and anomalous pixels.
We have further adapted AnomalyDINO to the batched zero-shot setting (see Appendix D). This adaptation is straightforward and relatively simple, especially compared to MuSc. Notably, we did not employ additional techniques such as ‘Re-scoring with Constrained Image-level Neighborhood’ or ‘Local Neighborhood Aggregation with Multiple Degrees’ (Li et al., 2024b). The results obtained, as detailed in Table 4, are therefore quite satisfactory.
We conclude that AnomalyDINO—despite its simplicity—rivals other methods in all settings and even comes out ahead in most of the settings. Within the class of training-free models, it is the clear winner essentially across the board. These results do not only demonstrate the virtues of AnomalyDINO itself but, more generally, highlight the merits of strong visual features as compared to highly engineered architectures. We provide further qualitative results in Figures 3 and 4, and discuss the limitations (such as detecting semantic anomalies) and specific failure cases of our approach in Appendix B.

### v2

@@ 3.1 Anomaly Detection via Patch (Dis-) Similarities
Let us briefly review the idea of the memory bank approach. We assume to have access to a suitable feature extractor $f:\mathcal{X}\rightarrow\mathcal{F}^{n}$ that maps each image $\mathbf{x}\in\mathcal{X}$ to a tuple of patch-features $f(\mathbf{x})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$, where $\mathcal{X}$ denotes some space of images and $\mathcal{F}$ the feature space. Note that $n$ depends on the image resolution and the patch size. Given $k\geq 1$ nominal reference sample(s) $X_{\text{ref}}:=\bigl{\{}\mathbf{x}^{(i)}\mid\,i\in[k]\bigr{\}}$ (with shorthand $[k]:=\{1,\dots,k\}$), we collect the nominal patch features and store them in a memory bank
@@
$\begin{split}\mathcal{M}:=\bigcup_{\mathbf{x}^{(i)}\in X_{\operatorname{ref}}}\bigl{\{}\mathbf{p}_{j}^{(i)}\mid f(\mathbf{x}^{(i)})=(\mathbf{p}_{1}^{(i)},\dots,\mathbf{p}_{n}^{(i)}),\,j\in[n]\bigr{\}}\enspace.\end{split}$ (1)
@@
To score a test sample $\mathbf{x}_{\mathrm{test}}$, we collect the extracted patch representations $f(\mathbf{x}_{\mathrm{test}})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n})$ and then check how well they comply with $\mathcal{M}$. To do so, we leverage a nearest-neighbor approach to find the distance to the closest reference patch for a given test patch $\mathbf{p}\in\mathcal{F}$
@@
Throughout this paper, we define $q$ as the average distance of the 1% most anomalous patches, i.e, $q(\mathcal{D}):=\mathrm{mean}(H_{0.01}(\mathcal{D}))$ with $H_{0.01}(\mathcal{D})$ containing the 1% highest values in $\mathcal{D}$. The statistic $q$ can be understood as an empirical estimate of the tail value at risk for the 99% quantile [26] and turns out suitable for a wide range of settings because it balances two desirable properties: First, we want $s(\mathbf{x}_{\mathrm{test}})$ to depend on the highest patch distances, as they may provide the strongest anomaly signal. Similarly, we want a certain degree of robustness against a singular high patch distance (in particular in few-shot scenarios where $\mathcal{M}$ is sparsely populated). However, one could also replace $q$ with a different statistic to cater to special cases. If anomalies are expected to cover larger parts of the image, for example, a high percentile of the patch distances might be a suitable choice. If, on the contrary, anomalies may occur only locally such that only very few patches/or a single patch might be affected, the score $q$ should be sensitive to the highest patch distances. Frequently, the maximum pixel-wise anomaly score is considered. Such an anomaly score on pixel-level is usually obtained by upsampling to full image resolution and applying some smoothing operation. Following [33], we utilize bilinear upsampling and Gaussian smoothing ($\sigma=4.0$) to turn the patch distances into pixel-level anomaly scores for localization of potential defects. Examples of the resulting anomaly maps are visualized in Figure 1 and Appendix A (Figures 4 and 5).
@@ 3.2 Enriching the Memory Bank & Filtering Relevant Patches
In the few-shot anomaly detection setting, the primary challenge is to effectively capture the concept of normality from the limited set of nominal samples. A useful strategy involves applying simple augmentations, like rotations (following the insights in [41]), to enhance the variability of nominal patch features. This is essential because the variability at test time is likely to be significantly greater than in the reference data, $X_{\mathrm{ref}}$.
To avoid irrelevant areas of the test image from leading to falsely high anomaly scores, we propose masking the object of interest. This approach mitigates the risk of false positives, particularly in the few-shot regime where the limited reference samples may not adequately capture the natural variations in the background, as exemplified in Figure 8, Appendix C.
It is important to note that the appropriate preprocessing technique—whether to apply masking and/or augmentations—should depend on the specific characteristics of the object of interest. For a more detailed discussion on the challenges and considerations in designing an effective preprocessing pipeline, see Section C.1.
@@
Masking, i.e., delineating the primary object(s) in an image from its background, can help to reduce false-positive predictions. To minimize the overhead of the proposed pipeline, we utilize DINOv2 also for masking. This is achieved by thresholding the first PCA component of the patch features [29]. We observe empirically that this sometimes produces erroneous masks. Frequently, such failure cases occur for close-up shots, where the objects of interest account for $\gtrsim 50\%$ of patches. To address this issue, we check if the PCA-based mask accurately captures the object in the first reference sample and apply the mask accordingly, which gives rise to the ‘masking test’ in Figure 2. This test is performed only once per object as the procedure yields very consistent outputs. In addition, we utilize dilation and morphological closing to eliminate small holes and gaps within the predicted masks. See Figures 12 and 13 for examples of this masking procedure, Table 8 for the outcomes of the masking test per object, and Figure 8 for a visualization of the benefits of masking in the presence of background noise (all in Appendix C). In general, we do not mask textures (e.g., ‘Wood’ or ‘Tile’ in MVTec-AD).
Figure 2: Masking test on MVTec-AD. For ‘Capsule’ and ‘Hazelnut’ the masking works successfully (top row), while for ‘Cable’ and ‘Transistor’ (bottom row) some areas are incorrectly predicted as background that should belong to the object of interest (highlighted in red). See App. C.1 for the outcomes per object.
@@
Rotating the reference sample may improve the detection performance by better resembling the variations within the concept of normality captured in $\mathcal{M}$. Consider, e.g., the ‘Screw’ in MVTec-AD as depicted in Figure 1 for which rotation-invariant features are desirable. On the other hand, rotation can also be detrimental in cases where rotations of (parts of) the object of interest can be considered anomalies themselves (see Figure 11 for an example).
We consider two different settings. In the ‘agnostic’ case, which we focus on in the main paper, we always augment the reference sample with rotations. We also consider the case where we know about the samples’ potential rotations (‘informed’). The ‘informed’ case is sensible as the data collection process can usually be controlled in an industrial/medical setting (or test images can be aligned), and may lead to lower inference time. See Section C.1 for the effect of masking and rotation, and a comparison of ‘informed’ vs. ‘agnostic’.
@@ 4 Experiments
Using DINOv2111Code/weights of DINOv2 [29] are available under Apache 2.0 license. as backbone allows us to choose from different distillation sizes, which range from small (ViT-S, $21\times 10^{6}$ parameters) to giant (ViT-G, $1.1\times 10^{9}$ parameters). To prioritize low latency, we take the smallest model as our default (and denote our default pipeline AnomalyDINO-S, accordingly) and evaluate two input resolutions, 448 and 672 pixels (smaller edge). As discussed above, we utilize the ‘agnostic’ preprocessing by default (see Appendix C).
@@
For the experiments, we consider MVTec-AD [^2] The MVTec-AD dataset [2] is available under CC BY-NC-SA 4.0. and VisA, [^3] The VisA dataset [48] is released under the CC BY 4.0 license. two datasets with high-resolution images for industrial anomaly detection. MVTec-AD consists of fifteen categories, depicting single objects or textures, and up to eight anomaly types per category. VisA has twelve categories and comes with diverse types of anomalies condensed into one anomaly class ‘bad’. Some categories in VisA can be considered more challenging as multiple objects and complex structures are present.
@@
We assess the image-level detection and the pixel-level segmentation performance with three metrics each, following [18, 7, 23]. For evaluating the detection performance, we measure the area under the receiver-operator curve (AUROC), the F1-score at the optimal threshold (F1-max), and the average precision (AP) using the respective image-level anomaly scores. We quantify the anomaly segmentation performance using the AUROC, F1-max, and the per-region overlap (PRO, [2]) of the segmentation using the pixel-wise anomaly scores. Note that due to a high imbalance of nominal and anomalous pixels—for MVTec-AD we have 97.26% nominal pixel, for VisA even 99.45% [23]—it is not recommended to assess performance solely on segmentation AUROC [31]. We repeat each experiment three times and report mean and standard deviation. [^4] The randomness stems from the choice of reference samples $X_{\mathrm{ref}}$. For straightforward reproducibility, we simply set $X_{\mathrm{ref}}$ to the first $k$, second $k$ and third $k$ train samples for the three different runs, respectively.
Baselines
We compare AnomalyDINO with a range of modern zero- and few-shot AD models, e.g., SPADE [8], PaDiM [10], PatchCore [33], WinCLIP+ [18], and APRIL-GAN [7]. It is important to note that ACR [21] and MuSc [23] consider the batched zero-shot setting (see Table 1), thus, covering a different setting than AnomalyDINO. We adapt AnomalyDINO to this setting in Appendix D. Moreover, APRIL-GAN and AnomalyCLIP, require training on a related dataset, which is in contrast to our proposed training-free method. In Tables 2 and 3, results reported by [18] are indicated by , the result of the WinCLIP re-implementation by [20] (where J. Jeong is also coauthor) by . All other results are taken from the original publication. For ADP [20] we report the (usually slightly better) variant ADP (with access to the class label). For GraphCore no standard deviations are reported [41].
4.1 Few-Shot Anomaly Detection and Anomaly Segmentation
Results
We summarize the results for few-shot anomaly detection on MVTec-AD and VisA in Table 2 and Table 3, respectively.
@@ Results
Method Classification Segmentation
@@
AnomalyDINO-S (448) 96.5$\pm$0.4 96.0$\pm$0.2 98.1$\pm$0.3 96.3$\pm$0.1 57.9$\pm$0.8 91.7$\pm$0.1
AnomalyDINO-S (672) 96.6$\pm$0.4 95.8$\pm$0.5 98.2$\pm$0.2 96.8$\pm$0.1 60.2$\pm$1.1 92.7$\pm$0.1
2-shot
SPADE 81.0$\pm$2.0 90.3$\pm$0.8 90.6$\pm$0.8 91.2$\pm$0.4 42.4$\pm$1.0 83.9$\pm$0.7
@@
AnomalyDINO-S (448) 96.7$\pm$0.8 96.5$\pm$0.4 98.1$\pm$0.7 96.5$\pm$0.2 58.5$\pm$0.5 92.0$\pm$0.2
AnomalyDINO-S (672) 96.9$\pm$0.7 96.1$\pm$0.3 98.2$\pm$0.5 97.0$\pm$0.2 61.0$\pm$0.5 93.1$\pm$0.2
4-shot
SPADE 84.8$\pm$2.5 91.5$\pm$0.9 92.5$\pm$1.2 92.7$\pm$0.3 46.2$\pm$1.3 87.0$\pm$0.5
@@
AnomalyDINO-S (448) 98.3$\pm$0.1 97.7$\pm$0.2 99.3$\pm$0.0 97.1$\pm$0.1 60.0$\pm$0.1 92.9$\pm$0.1
AnomalyDINO-S (672) 98.4$\pm$0.1 97.6$\pm$0.1 99.3$\pm$0.0 97.5$\pm$0.0 62.7$\pm$0.1 94.0$\pm$0.1
Regarding MVTec-AD, our method achieves state-of-the-art $k$-shot detection performance across all $k\in{1,2,4,8,16}$ for every reported metric, outperforming approaches that require additional training data sets (such as ADP and APRIL-GAN). The method also demonstrates superior anomaly localization, with results showing that while detection performance remains comparable across different resolutions (448 vs. 672), a higher resolution enhances localization performance. Furthermore, we observe clear improvements as the number of samples increases.
@@
Method Classification Segmentation
@@
AnomalyDINO-S (448) 93.8$\pm$0.1 89.9$\pm$0.3 94.2$\pm$0.3 98.3$\pm$0.0 48.6$\pm$0.3 93.8$\pm$0.2
AnomalyDINO-S (672) 94.8$\pm$0.2 90.9$\pm$0.2 95.3$\pm$0.3 98.5$\pm$0.0 52.5$\pm$0.5 95.3$\pm$0.2
In terms of anomaly detection in the VisA benchmark (Table 3), APRIL-GAN demonstrates the strongest performance for $k\in\{1,2\}$. Nonetheless, AnomalyDINO consistently achieves second-best results for $k\in\{1,2\}$, comparable results in the 4-shot setting, and sets new state-of-the-art for $k\in\{8,16\}$. This can be attributed to AnomalyDINO’s ability to benefit more from a richer memory bank $\mathcal{M}$ than APRIL-GAN. We hypothesize that meta-learning exerts a greater influence on APRIL-GAN (i.e., training on MVTec-AD, when testing on VisA, and vice-versa) compared to learning from the nominal features of the given reference samples. Note also that AnomalyDINO outperforms all other training-free approaches.
Comparing the segmentation performance on VisA, Table 3 reveals a clear picture: AnomalyDINO consistently shows the strongest localization performance in all metrics considered. While AnomalyDINO-S (448) already demonstrates strong performance, the advantages of using a higher resolution (672) become more evident on the VisA dataset. We attribute this fact to smaller anomalous regions (for which smaller effective patch sizes are beneficial) and more complex scenes (compared to MVTec-AD).
We have further adapted AnomalyDINO to the batched zero-shot setting (see Appendix D). This adaptation is straightforward and relatively simple, especially compared to MuSc. Notably, we did not employ additional techniques such as ‘Re-scoring with Constrained Image-level Neighborhood’ or ‘Local Neighborhood Aggregation with Multiple Degrees’ [23]. The results obtained, as detailed in Table 4, are therefore quite satisfactory.
We conclude that AnomalyDINO—despite its simplicity—rivals other methods in all settings and even comes out ahead in most of the settings (e.g., significantly reducing the gap between few-shot and full-shot methods for MVTec-AD). Within the class of training-free models, it is the clear winner essentially across the board. These results do not only demonstrate the virtues of AnomalyDINO itself but, more generally, highlight the merits of strong visual features as compared to highly engineered architectures. We provide further qualitative results in Appendix A, and discuss the limitations (such as detecting semantic anomalies) and specific failure cases of our approach in Appendix B.
Table 4: Detection results for other settings. All results are AUROC values (in %).
@@
AnomalyDINO-S (448) 93.0 89.7
AnomalyDINO-S (672) 94.2 90.7